# Chaos-based reinforcement learning with TD3

### v1

@@ 3.3 TD3-CBRL
The following modifications are made to the TD3 algorithm to adapt to the CBRL approach. We eliminated the random number exploration noise $\epsilon^{a}$ and $\epsilon^{t}$. Instead of exploration by random numbers, we rely on exploration driven spontaneously by chaotic dynamics and use an ESN with a larger spectral radius, as shown in Fig. 3(a), for the $\mu$ network. Here, $g=2.2$ unless otherwise mentioned. We also add the reservoir output to the experience stored in the replay buffer. That is, $e_{t}=(\bm{u}_{t},\bm{x}_{t},a_{t},r_{t},\bm{u}_{t+1},\bm{x}_{t+1})$ is stored into the replay buffer $\mathcal{B}$. The agent learns without the Back Propagation Through Time method using the stored $\bm{u}$ and $\bm{x}$ as state $s_{t}$. This method has been employed in several studies and efficiently trains deep reinforcement learning agents using ESN [59, 60]. TD3 uses random numbers to randomly sample past experiences from the replay buffer. Since this sampling has an exploration-like effect with the random numbers, the replay buffer and batch sizes are set to the same value of $64$ to eliminate this effect. Therefore, the training is based on the memory of the past $64$ steps of experience. The variant of CBRL that introduces the above modified TD3 algorithm is called TD3-CBRL in this study. In the experiment, we compare the three cases shown in Fig. 3 (a-c) to confirm that the ESN dynamics contributes to the exploration during learning.
Fig. 4: Goal Task. The blue cylinder represents the agent, whose position is denoted by $x^{\mathrm{A}}$ and $y^{\mathrm{A}}$. The field is surrounded by walls, and $x^{\mathrm{A}}$ and $y^{\mathrm{A}}$ cannot be out of the range from 00 to $20$. The agent can move horizontally and vertically within the environment, with a maximum distance of $1$ each. The yellow circle indicates the goal, and if the agent’s center is within this circle, the agent is rewarded and considered to have accomplished the task. The $4$ red dots indicate the initial positions. At the beginning of each episode, the agent starts at one of these coordinates, with a slight noise added to the position.
@@ 3.4 Goal task
This study uses a goal task to estimate the agent’s learning ability. Figure 4 shows the task outline. The task field is inside the range 00 to $20$ on the x-y plane. There are $4$ initial positions $(x,y)=(2,2),(2,18),(18,2),(18,18)$ in the environment. At the beginning of an episode, the agent is randomly placed at one of the initial positions, with Gaussian noise $\mathcal{N}(0,1)$ added to its position in the field. The agent obtains an input $\bm{u}$ based on its position $(x^{\mathrm{A}},y^{\mathrm{A}})$ and uses it to determine its action outputs $-1\leq a^{x}\leq 1$ and $-1\leq a^{y}\leq 1$. As a result of the action, the agent’s position is updated according to the following equation
@@
[…]
where $D_{\mathrm{G}}$ is the Euclidean distance between the center of the agent and the goal, and $L$ is the length of the diagonal of the task field. This task is episodic, and an episode ends when the agent either enters the goal circle of radius $2$ (i.e., $D_{\mathrm{G}}<2$) or fails to reach the goal within $200$ steps. The agent receives a reward of $r=1$ for reaching the goal, $r=-0.01$ for colliding with the wall surrounding the field, and $r=0$ for any other step.
@@ 4.1 Conditions
The reservoir size was set to $N_{x}=256$ and the recurrent connection probability was set to $p=0.1$. The input weight matrix $W_{\mathrm{in}}$ was sampled from a uniform distribution over $[-0.5,0.5]$. The critic network is a fully connected Multilayer Perceptron (MLP) with two hidden layers consisting of $32$ ReLU nodes, and the output neuron is a linear node. The initial weights were set using the default settings in PyTorch (version 1.13.1 in this study.). Both the critic $Q$ and the actor $\mu$ are trained using the Adam optimizer, with a learning rate of $0.0003$. We paused training every $N_{v}$ steps and verified the agent’s behavior starting from 8 initial positions $(2,2)$, $(2,10)$, $(2,18)$, $(10,2)$, $(10,18)$, $(18,2)$, $(18,10)$, $(18,18)$ and slightly different positions (shifted $0.002$ in the $x$ and $y$ axes, respectively). The replay buffer size and batch size were both set to 64. The discount rate $\gamma$ was set to 0.95. The time constant $\tau$ of the target network was set to 0.05. We set $\epsilon^{a}=0,\epsilon^{t}=0$, and the Delayed Policy Updates step to $d=2$.
@@ 4.2 Learning result
The TD3-CBRL agent learned the goal task in $20000$ steps. Figure 5(a) shows the learning curve resulting from the test conducted every $N_{v}=2000$ step. This figure shows that the average number of steps required by the agent to reach the goal decreases as the learning progresses. This result indicates that TD3-CBRL can successfully learn the goal task. Figure 6(a) shows the trajectories of the agent’s movement in the environment for each test trial. This figure shows that in the initial stage of learning (00 step), the agent acts exploratively in the environment. On the other hand, in the trajectories after $4000$ steps, the agent is moving toward the goal from each initial position, although it takes detours in some cases. We also see that the exploratory behavior subsides as the learning progresses. It is important to note that no external random noises for exploration were added to the agent’s action output during the learning process. This result indicates that the internal dynamics of the reservoir caused spontaneous exploratory behavior, and as learning progressed, such variability in the output was autonomously suppressed.
@@
Fig. 6: Agent trajectories during the test. Each colored trajectory represents the agent’s behavior from the $8$ initial test positions. Each graph in the subfigures shows the test results at 00, $4000$, $8000$, $12000$, $16000$, and $20000$ training steps. The random seed is the same as the one used for the representative results in Fig. 5. (a) shows the training results for TD3-CBRL. (b) shows the results of TD3 without exploration by random noises. (c) shows the results with regular TD3. (d) shows the behavior of the same agent as in (c) when the exploration by random noises is not stopped during the test.
As seen in the purple trajectory where the agent starts from the initial position $(10,18)$ in Fig. 6(a), the agent wandered to the left around the goal before reaching it at the $20000$ step, even though the agent had already learned to reach the goal in a straight path by the $12000$ step. This result seems to be caused by the chaoticity of the agent’s dynamics, which causes its behavior to change sensitively depending on the initial position and parameter changes. Figure 7 (a) shows the results of investigating the sensitivity of the agent’s behavior to variations of the initial positions in this case. This figure shows the trajectories of agents starting from slightly different initial positions. At the 00 step, the agent, starting from its original initial position, continues toward the wall and ends the episode without reaching the goal. On the other hand, although the agent starting from the slightly shifted position behaves like the agent starting from the original initial position initially, it leaves the original trajectory after a while and eventually reaches the goal. This result indicates that in the early phases of learning, a slight difference in the initial position can significantly change the agent’s behavior and even determine the task’s success or failure. At the $8000$ step, the two agents eventually reach the goal, but their paths diverge from the middle of the episode. The agent, starting from the original initial position, explores both the left and right sides before reaching the goal. On the other hand, the agent starting from a slightly different point did not explore the right side of the goal; instead, it went to the goal. Thus, the sensitivity to the initial condition of the TD3-CBRL agent allows it to exhibit both exploration driven by the chaoticity and exploitation of its previously learned behavior.
@@
Fig. 7: Sensitivity of agent trajectories for initial position. The blue and orange lines show the agent’s trajectory when the agent started from (10, 18) and from a slightly shifted initial position (10.002, 18.002), respectively. Each graph in the subfigure shows the test results at $0,4000,8000,12000,16000$, and $20000$ training steps.
@@ 4.3 Effects by presence or absence of exploration component
To confirm that the chaoticity of the model contributes to the agent’s exploration, we evaluated the learning performance when the model does not have chaotic dynamics. Specifically, instead of an actor-network with a chaotic reservoir, we used a Multilayer Perceptron (MLP) with one hidden layer consisting of 256 $\tanh$ neurons, as shown in Fig. 3 (b). In this case, the learning rate was readjusted to $0.00003$. Figure 5(b) shows the learning curve for the MLP case without exploration random noise. This figure shows that the number of steps to the goal did not decrease and that the agent failed to learn the goal task. Figure 6(b) and Fig. 7 (b) show the agent trajectory during the test phase and the sensitivity test under this condition. These figures show that the agent lacked exploratory behavior and failed to learn. This result appears to be due to the fact that the learning model did not have the dynamics to generate spontaneous activity, and thus, the exploration did not occur without random noises.
To clarify that the absence of exploration random noise is the direct cause of the learning failure, we verified the case where the MLP is trained with random noise for exploration in the same way as in the regular TD3, as shown in Fig. 3 (c). Figure 5(c) shows the learning curve under this condition. This figure shows that the number of steps to reach the goal decreased and that the agent succeeded in learning the task. Figure 6(c) shows the trajectories of the agent’s behavior in this validation. This figure shows that the agent can learn the behavior of moving toward the goal due to exploration with random numbers during the learning process. Note that during the test phase, adding random numbers to the action outputs is stopped. These results indicate that the presence or absence of exploration by random numbers has a significant influence on the success or failure of learning and that the TD3-CBRL agent successfully learns the goal task through exploration driven by the chaotic dynamics of the reservoir.
Comparing (a) and (c) in Fig. 6, the regular TD3 agent can go to the goal in a straighter path than the TD3-CBRL agent. However, this is due to an external intervention that removes the random noise during the test phase. Figure 6(d) shows the result when the agent in (c) acts with the exploration random noise during the test phase. If the exploration random noise is not eliminated, the agent cannot reach the goal in a straight path. On the other hand, the TD3-CBRL agent can autonomously reduce the variability driven by its chaoticity as its learning progresses, although there are still cases where it takes detours depending on its initial position. This is a beneficial property and a limitation of the CBRL agent.
@@
(b) Larger spectral radius case $(g=5)$.
@@
In the goal change task, the goal was initially placed at $G_{p}^{\mathrm{1}}=(15,10)$, and learning was performed for $20000$ steps. Then, at the $N_{\mathrm{c}}=20001$ step, the goal position was changed to $G_{p}^{\mathrm{2}}=(5,10)$, and learning continued in the new environment for another $30000$ steps. The test is conducted every $N_{v}=2000$ step. The learning curve under these conditions is shown in Fig. 8(a). This figure shows that when the goal position is changed, the number of steps required to reach the goal, which had decreased during learning in the initial environment, temporarily increases. However, as learning in the new environment progresses, the number of steps decreases again. This result indicates that the agent is adapting to the new environment.
@@
4.5 Learning performance and chaoticity
We investigated the effect of the chaoticity of the system on learning performance. The chaoticity of the reservoir network can be tuned by changing the spectral radius of the reservoir’s recurrent weight matrix using the parameter $g$. We changed the value of $g$ from 00 to $12$ in $0.2$ increments and conducted trials with 100 different random number seeds under each condition to obtain the learning success probability and the average number of steps to reach the goal. Here, success is defined as an event when the agent reaches the goal from all $16$ initial positions during the final test. The results are shown in Fig. 10(a). This figure shows that learning performance begins to improve as $g$ exceeds around $1$, and learning becomes successful in most cases when $g$ exceeds approximately $2$. This result indicates that the chaoticity of the system is essential for successful exploration and learning. The results also indicate that the probability of success remains high even when the value of $g$ is further increased. In general, there is an appropriate range for the parameter $g$ for time series processing in reservoir computing. It can be considered that since this study does not target tasks that require memory and the task is simple, the success probability of learning by TD3-CBRL remains stable even when $g$ is larger than the typical critical value.
@@ 4.5 Learning performance and chaoticity
Fig. 10: Learning performance with varying spectral radius. The blue line shows the learning success probability, indicated on the left vertical axis. The red line shows the average number of steps required to reach the goal, indicated on the right vertical axis. The horizontal axis shows the value of the spectral radius. (a) shows the learning results for the goal task. (b) shows the learning results of the goal change task. (c) shows the results of learning over a long period of 18,000 steps after the goal change.
@@
We examined how the flexibility of switching between exploration and exploitation was affected by the value of $g$. The value of $g$ was changed from 00 to $12$ in $0.2$ increments, and a goal change task was learned under $100$ different random number seeds to ensure statistical validity. The task settings are the same as in Section 4.4 except for $g$. The results are shown in Fig. 10(b). This figure shows that performance on the goal change task decreases with an extremely large $g$. This result suggests that when $g$ is extremely large, the balance between exploration and exploitation is lost, and the flexibility to deal with environmental changes is reduced. These results indicate that choosing an appropriate $g$ value is still essential in CBRL using reservoir networks.
@@
We experimented with long-term re-learning to investigate whether a larger $g$ decreases the model’s re-learning ability or increases the number of steps required for re-learning. Specifically, after changing the goal position at the $N_{\mathrm{c}}=20001$ step, the agent learned for $180000$ steps in the new environment. The results are shown in Fig. 10(c). This figure shows that long-term learning mitigated the decrease in learning performance when $g$ is large. This result indicates that a model with a larger $g$ and stronger chaoticity requires more re-learning steps, consequently making re-learning more difficult.
4.6 Readout weights and chaoticity
To clarify the reason for the increase in steps required for re-learning when $g$ is extremely large, we investigated how the acquired readout weights vary with the spectral radius $g$, which adjusts the chaoticity of the reservoir. The inputs for the readout are the reservoir output and the agent’s state in the environment. They are received through 256 and 5 weights, respectively. Figure 11 shows the readout weights after training for spectral radius $g$ of $2.2$ and $5$. The blue dots show the weights from the reservoir, and the red dots show the bypass weights. Comparing these figures, we can see that the bypass weights are larger for the case of $g=5$ than for the case of $g=2.2$, and the weights from the reservoir are relatively smaller than the bypass weights. Figure 12 shows how the average of the absolute values of the weights from the reservoir and the bypass weights change when the spectral radius is varied. This figure shows that the bypass weights increase as the spectral radius increases. This result suggests that the agent ignores the reservoir outputs and focuses more on the state inputs directly given from the environment.
@@ 4.6 Readout weights and chaoticity
(a) $g=2.2$
(b) $g=5$
Fig. 13: Changes in the weights of the bypass. The vertical axis shows the weights, and the horizontal axis shows the number of training steps. (a) and (b) show the results at $g=2.2$ and $g=5$, respectively. The upper and lower figures show the weights of the readout unit in the $x$-axis and $y$-axis directions, respectively.
We have validated the goal change task with $g=5$. Figure 8(b) shows the results of this validation. Comparing this with the result under the setting of $g=2.2$ shown in Fig. 8(a), re-learning convergence was slower when $g=5$. Figure 9(b) shows the agent’s behavior in the environment at $g=5$. Comparing Fig. 9(a) and (b), we can see that in the case of $g=5$, the agent continues its exploratory behavior for more steps and takes more steps to shift to an exploitation mode. This slow re-learning seems to be because more updates are required to modify the larger bypass weights to adapt them to the new environment when the spectral radius is large. To clarify this, we analyzed the change in the bypass weights in the re-learning task, and Fig. 13 shows the results. Comparing the two results, the bypass weight change in the case of $g=5$ is significantly larger than in the case of $g=2.2$. In the case of $g=5$, the bypass weights increased to attend to state inputs, ignoring the reservoir’s chaoticity, and thus, more steps are required. These results confirm that the inflated bypass weights inhibit flexible switching between exploration and exploitation. These results indicate that an excessive spectral radius and agent chaoticity impair the ability of TD3-CBRL agents to properly balance exploration and exploitation properly.

### v2

@@ 3.3 TD3-CBRL
The following modifications are made to the TD3 algorithm to adapt to the CBRL approach. We eliminated the random number exploration noise $\epsilon^{a}$ and $\epsilon^{t}$. Instead of exploration by random numbers, we rely on exploration driven spontaneously by chaotic dynamics and use an ESN with a larger spectral radius, as shown in Fig. 3(a), for the $\mu$ network. Here, $g=2.2$ unless otherwise mentioned. We also add the reservoir state to the experience stored in the replay buffer. That is, $e_{t}=(\bm{u}_{t},\bm{x}_{t},a_{t},r_{t},\bm{u}_{t+1},\bm{x}_{t+1})$ is stored into the replay buffer $\mathcal{B}$. The agent learns without the Back Propagation Through Time method using the stored $\bm{u}$ and $\bm{x}$ as state $s_{t}$. This method has been employed in several studies and efficiently trains deep reinforcement learning agents using ESN [50, 51]. The variant of CBRL that introduces the above modified TD3 algorithm is called TD3-CBRL in this study. In the experiment, we compare the three cases shown in Fig. 3 (a-c) to confirm that the ESN dynamics contributes to the exploration during learning.
Fig. 4: Goal task. The blue cylinder represents the agent, whose position is denoted by $x^{\mathrm{A}}$ and $y^{\mathrm{A}}$. The field is surrounded by walls, and $x^{\mathrm{A}}$ and $y^{\mathrm{A}}$ cannot be out of the range from 0 to $20$. The agent can move horizontally and vertically within the environment, with a maximum distance of 11 each. The yellow circle indicates the goal, and if the agent’s center is within this circle, the agent is rewarded and considered to have accomplished the task. The 44 red dots indicate the initial positions. At the beginning of each episode, the agent starts at one of these coordinates, with a slight noise added to the position.
@@ 3.4 Goal task
This study uses a goal task to estimate the agent’s learning ability. Figure 4 shows the task outline. The task field is inside the range 0 to $20$ on the x-y plane. There are $4$ initial positions $(x,y)=(2,2),(2,18),(18,2),(18,18)$ in the environment. At the beginning of an episode, the agent is randomly placed at one of the initial positions, with Gaussian noise $\mathcal{N}(0,1)$ added to its position in the field. The agent obtains an input $\bm{u}$ based on its position $(x^{\mathrm{A}},y^{\mathrm{A}})$ and uses it to determine its action outputs $1\leq a^{x}\leq 1$ and $1\leq a^{y}\leq 1$. As a result of the action, the agent’s position is updated according to the following equation
@@
$\bm{u}=\left[\frac{x^{\mathrm{A}}}{20}\ \ 1-\frac{x^{\mathrm{A}}}{20}\ \ \frac{y^{\mathrm{A}}}{20}\ \ 1-\frac{y^{\mathrm{A}}}{20}\ \ 1-\frac{D_{\mathrm{G}}}{2L}\right]^{\top},$ (14)
where $D_{\mathrm{G}}$ is the Euclidean distance between the center of the agent and the goal, and $L$ is the length of the diagonal of the task field. This task is episodic, and an episode ends when the agent either enters the goal circle of radius $2$ (i.e., $D_{\mathrm{G}}<2$) or fails to reach the goal within $200$ steps. The agent receives a reward of $r=1$ for reaching the goal, r=0.01r=-0.01 for colliding with the wall surrounding the field, and $r=0$ for any other step.
@@ 4.1 Conditions
The reservoir size was set to $N_{x}=256$ and the recurrent connection probability was set to $p=0.1$. The input weight matrix $W_{\mathrm{in}}$ was sampled from a uniform distribution over [0.5,0.5][-0.5,0.5]. The critic network $Q$ is a fully connected Multilayer Perceptron (MLP) with two hidden layers consisting of $32$ ReLU nodes, and the output neuron is a linear node. The initial weights were set using the default settings in PyTorch (version 1.13.1 in this study.). Both the critic network $Q$ and the actor network $\mu$ are trained using the Adam optimizer, with a learning rate of $5\times 10^{-4}$. We paused training every $N_{v}$ steps and verified the agent’s behavior starting from 8 initial positions $(2,2)$, $(2,10)$, $(2,18)$, $(10,2)$, $(10,18)$, $(18,2)$, $(18,10)$, $(18,18)$ and slightly different positions (shifted $0.002$ in the $x$ and $y$ axes, respectively). The replay buffer size was set to $10^{6}$ and the batch size was set to 64. The discount factor $\gamma$ was set to 0.95. The time constant $\tau$ of the target network was set to 0.05. We set $\epsilon^{a}=0,\epsilon^{t}=0$, and the Delayed Policy Updates step to $d=2$.
@@ 4.2 Learning result
The TD3-CBRL agent learned the goal task in $20000$ steps. Figure 5(a) shows the learning curve resulting from the test conducted every $N_{v}=2000$ step. This figure shows that the average number of steps required by the agent to reach the goal decreases as the learning progresses. This result indicates that TD3-CBRL can successfully learn the goal task. Figure 6(a) shows the trajectories of the agent’s movement in the environment for each test trial. This figure shows that in the initial stage of learning (0 step), the agent acts exploratively in the environment. On the other hand, in the trajectories after $4000$ steps, the agent is moving toward the goal from each initial position. We also see that the exploratory behavior subsides as the learning progresses. It is important to note that no external random noises for exploration were added to the agent’s action output during the learning process. This result indicates that the internal dynamics of the reservoir caused spontaneous exploratory behavior, and as learning progressed, such variability in the output was autonomously suppressed.
@@
Fig. 6: Agent trajectories during the test. Each colored trajectory represents the agent’s behavior from the $8$ initial test positions. Each graph in the subfigures shows the test results at 0, $4000$, $8000$, $12000$, $16000$, and $20000$ training steps. The random seed is the same as the one used for the representative results in Fig. 5. (a) shows the training results for TD3-CBRL. (b) shows the results of TD3 without exploration by random noises. (c) shows the results with regular TD3. (d) shows the behavior of the same agent as in (c) when the exploration by random noises is not stopped during the test.
Figure 7 (a) shows the results of investigating the sensitivity of the agent’s behavior to variations of the initial positions in this case. This figure shows the trajectories of agents starting from slightly different initial positions. At the 0 step, the agent, starting from its original initial position, continues toward the wall and ends the episode without reaching the goal. On the other hand, although the agent starting from the slightly shifted position behaves like the agent starting from the original initial position initially, it leaves the original trajectory after a while and eventually reaches the goal. This result indicates that in the early phases of learning, a slight difference in the initial position can significantly change the agent’s behavior and even determine the task’s success or failure. At the $20000$ steps, the trajectories starting from two different initial positions reach the goal before they diverge.
@@
Fig. 7: Sensitivity of agent trajectories for initial position. The blue and orange lines show the agent’s trajectory when the agent started from (10, 18) and from a slightly shifted initial position (10.002, 18.002), respectively. Each graph in the subfigure shows the test results at 0 and $20000$ training steps.
@@ 4.3 Effects by presence or absence of exploration component
To confirm that the chaoticity of the model contributes to the agent’s exploration, we evaluated the learning performance when the model does not have chaotic dynamics. Specifically, instead of an actor network with a chaotic reservoir, we used a Multilayer Perceptron (MLP) with one hidden layer consisting of 256 $\tanh$ neurons, as shown in Fig. 3 (b). In this case, the learning rate of the actor network was readjusted to $1.6\times 10^{-5}$. Figure 5(b) shows the learning curve for the MLP case without exploration random noise. This figure shows that the number of steps to the goal scarcely decrease and that the agent failed to learn the goal task. Figure 6(b) and Fig. 7 (b) show the agent trajectory during the test phase and the sensitivity test under this condition. These figures show that the agent lacked exploratory behavior and failed to learn. This result appears to be due to the fact that the learning model did not have the dynamics to generate spontaneous activity, and thus, the exploration did not occur without random noises.
To clarify that the absence of exploration random noise is the direct cause of the learning failure, we verified the case where the MLP is trained with random noise for exploration in the same way as in the regular TD3, as shown in Fig. 3 (c). Here, $\epsilon^{a}$ is sampled from $\mathcal{N}(0,0.5^{2})$, and $\epsilon^{t}$ is sampled from $\mathcal{N}(0,0.2^{2})$ clipped to the range from -1 to 1. Figure 5(c) shows the learning curve under this condition. This figure shows that the number of steps to reach the goal decreased and that the agent succeeded in learning the task. Figure 6(c) shows the trajectories of the agent’s behavior in this validation. This figure shows that the agent can learn the behavior of moving toward the goal due to exploration with random numbers during the learning process. Note that during the test phase, adding random numbers to the action outputs is stopped. These results indicate that the presence or absence of exploration by random numbers has a significant influence on the success or failure of learning and that the TD3-CBRL agent successfully learns the goal task through exploration driven by the chaotic dynamics of the reservoir.
Comparing (a) and (c) in Fig. 6, the regular TD3 agent can go to the goal in a straighter path than the TD3-CBRL agent. However, this is due to an external intervention that removes the random noise during the test phase. Figure 6(d) shows the result when the agent in (c) acts with the exploration random noise during the test phase. If the exploration random noise is not eliminated, the agent cannot reach the goal in a straight path. On the other hand, the TD3-CBRL agent can autonomously reduce the variability driven by its chaoticity as its learning progresses.
@@
(b) Regular TD3.
(c) Larger spectral radius case $(g=5)$.
@@
In the goal change task, the goal was initially placed at $G_{p}^{\mathrm{1}}=(15,10)$, and learning was performed for $20000$ steps. Then, at the $N_{\mathrm{c}}=20001$ steps, the goal position was changed to $G_{p}^{\mathrm{2}}=(5,10)$, and learning continued in the new environment for another $30000$ steps. The test is conducted every $N_{v}=5000$ steps. The learning curve under these conditions is shown in Fig. 8(a). This figure shows that when the goal position is changed, the number of steps required to reach the goal, which had decreased during learning in the initial environment, temporarily increases. However, as learning in the new environment progresses, the number of steps decreases again. This result indicates that the agent is adapting to the new environment.
@@
Figure 8 (b) shows the results of applying the goal change task to the regular TD3 agent. This figure demonstrates that the regular TD3 agent fails to re-learn the new goal. While there is a possibility of task dependence, it suggests that TD3-CBRL agents exhibit greater flexibility in adapting to new environments compared to regular TD3 agents in goal change task. Two hypotheses can be proposed to explain this advantage of the CBRL agents. First, the short-term memory capability of the reservoir allows the agent to distinguish between the state after and before reaching the initial goal area. Second, the exploratory behavior driven by the chaotic dynamics of the reservoir is more effective than exploration based on simply adding random noise to the actions.
@@
(d) Goal change task with a replay buffer size of 64.
(e) Goal task with expanded task field.
(f) Goal task with observation noises.
@@
Fig. 10: Learning performance with varying spectral radius. The blue line shows the successful learning probability, indicated on the left vertical axis. The red line shows the average number of steps required to reach the goal, indicated on the right vertical axis. The horizontal axis shows the value of the spectral radius. (a) shows the results for the goal task. (b) shows the results of the goal change task. (c) shows the results of learning over a long period of 180000 steps after the goal change. (d) shows the results with a replay buffer size of $64$. (e) shows the results for the goal task whose task field was expanded from the range $[0,20]$ to $[0,100]$ in the x-y plane. (f) shows the results for the goal task with observation noises.
4.5 Learning performance and chaoticity
We investigated the effect of the chaoticity of the system on learning performance. The chaoticity of the reservoir network can be tuned by changing the spectral radius of the reservoir’s recurrent weight matrix using the parameter $g$. We changed the value of $g$ from 0 to $12$ in $0.2$ increments and conducted trials with 100 different random number seeds under each condition to obtain the successful learning probability and the average number of steps to reach the goal. Here, success is defined as an event when the agent reaches the goal from all $16$ initial positions during the final test. The results are shown in Fig. 10(a). This figure shows that learning performance begins to improve as $g$ exceeds around $1$, and learning becomes successful in most cases when $g$ exceeds approximately $2$. This result indicates that the chaoticity of the system is essential for successful exploration and learning. The results also indicate that the probability of success remains high even when the value of $g$ is further increased. In general, there is an appropriate range for the parameter $g$ for time series processing in reservoir computing. It can be considered that since this study does not target tasks that require memory and the task is simple, the success probability of learning by TD3-CBRL remains stable even when $g$ is larger than the typical critical value.
@@ 4.5 Learning performance and chaoticity
We conducted the same experiment with a goal change task to examine how the flexibility of switching between exploration and exploitation was affected by the value of $g$. The task settings are the same as in Section 4.4 except for $g$. The results are shown in Fig. 10(b). This figure shows that performance on the goal change task decreases with an extremely large $g$. These results indicate that choosing an appropriate $g$ value is still essential in CBRL using reservoir networks.
@@
We experimented with long-term re-learning to investigate whether a larger $g$ makes the agent unable to re-learn or increases the required steps. Specifically, after changing the goal position at the $N_{\mathrm{c}}=20001$ steps, the agent learned for $180000$ steps in the new environment. The results are shown in Fig. 10(c). This figure shows that long-term learning mitigated the decrease in learning performance when $g$ is large. This result indicates that a model with a larger $g$ and stronger chaoticity requires more re-learning steps, consequently making re-learning more difficult.
@@
We investigated how the acquired readout weights vary with the spectral radius $g$, which adjusts the chaoticity of the reservoir. The inputs for the readout are the reservoir state and the agent’s state in the environment. They are received through 256 and 5 weights, respectively. Figure 11 shows the readout weights after training for spectral radius $g$ of $2.2$ and $5$. The blue dots show the weights from the reservoir, and the red dots show the bypass weights. Comparing these figures, we can see that the bypass weights are larger for the case of $g=5$ than for the case of $g=2.2$, and the weights from the reservoir are relatively smaller than the bypass weights. Figure 12 shows how the average of the absolute values of the weights from the reservoir and the bypass weights change when the spectral radius is varied. This figure shows that the bypass weights increase as the spectral radius increases. This result suggests that as $g$ becomes excessively large, the agent tends to ignore the reservoir states and focus more on the inputs provided directly from the environment. This indicates that the states of a strongly chaotic reservoir lose their value as an input for the agent to accomplish the task.
@@
We have validated the goal change task with $g=5$. Figure 8(c) shows the results of this validation. Comparing this with the result under the setting of $g=2.2$ shown in Fig. 8(a), re-learning convergence was slower when $g=5$. Figure 9(b) shows the agent’s behavior in the environment at $g=5$. Comparing Figs. 9(a) and (b), we can see that in the case of $g=5$, the agent continues its exploratory behavior for more steps and takes more steps to shift to an exploitation mode. This slow re-learning seems to be because the strongly chaotic reservoir states stored in the replay buffer before the goal change negatively affected learning. To verify this, we set the replay buffer size to $64$, the same as the batch size, and trained relying only on recent experiences. Figure 10(d) shows the learning performance for different spectral radiuses under this setting. This figure shows that the required steps for re-learning did not increase even with large $g$, although the learning performance tends to decrease. This result suggests that excessive spectral radiuses complicate the reservoir state and reduce its correlation with the input from the environment. This makes the experience stored in the replay buffer before the goal change worthless for learning, thereby increasing the number of steps required for re-learning. Furthermore, excessively large $g$ hinders the TD3-CBRL agent’s ability to properly balance exploration and exploitation.
Excessive chaoticity in a reservoir can cause its states to diverge or fluctuate erratically. Such unstable dynamics poses a concern as they degrade the agent’s learning performance, especially in tasks that unfold over long time horizons. Therefore, to investigate the impact of the spectral radius $g$ on learning performance in a more extensive task field that requires long time steps to reach the goal, we experimented with the goal task whose field was expanded from the range $[0,20]$ to $[0,100]$ in the x-y plane. The result, shown in Fig. 10(e), demonstrates a marked degradation in performance at larger $g$ values. Consequently, the range for successful learning was restricted to an extremely narrow range centered around $g=1.8$. This result suggests that excessive chaoticity may hinder the formation of a global exploration in an extensive state space.
The sensitivity to initial conditions of the chaotic reservoir can cause the system to react sensitively to observation noise, potentially leading to performance instability. Therefore, to evaluate the robustness of TD3-CBRL against observation noise, we investigated its performance on the goal task when Gaussian noise $\mathcal{N}(0,10^{-2})$ was added to the observation $\bm{u}$ in the test phase. The results are shown in Fig. 10(f). This figure shows that, overall, no significant performance degradation was observed compared to the case without noise (Fig. 10(a)). This result suggests that the TD3-CBRL agent possesses a certain degree of robustness against observation noise under the experimental conditions of a simple goal task.